OpenStudy (anonymous):

What is the area of a circle with a circumfrence of 33 units?

OpenStudy (anonymous):

circumfrence of a circle=2pir=33units r=(33/2pi) =5.25 area of the circle=pir^2=pi*(5.25)^2=86.54625